# On some practical challenges of conformal prediction

### v1

@@ Abstract
Conformal prediction is a model-free machine learning method for creating prediction regions with a guaranteed coverage probability level. However, a data scientist often faces three challenges in practice: (i) the determination of a conformal prediction region is only approximate, jeopardizing the finite-sample validity of prediction, (ii) the computation required could be prohibitively expensive, and (iii) the shape of a conformal prediction region is hard to control. This article offers new insights into the relationship among the monotonicity of the non-conformity measure, the monotonicity of the plausibility function, and the exact determination of a conformal prediction region. Based on these new insights, we propose a simple strategy to alleviate the three challenges simultaneously.
Keywords and phrases: Data science; exact determination of conformal prediction regions; explainable machine learning; finite-sample validity.
@@ 1 Introduction
Suppose $Z_{1}=(X_{1},Y_{1}),Z_{2}=(X_{2},Y_{2}),\ldots$ is a sequence of exchangeable random vectors, where $X_{i}\in\mathbb{R}^{p}$ for $p\geq 1$, $Y_{i}\in\mathbb{R}$, and each $Z_{i}$ follows a distribution $\mathsf{P}$. Our goal is to perform an interval prediction of the next response $Y_{n+1}$ at a randomly sampled feature $X_{n+1}$, based on past observations of $Z^{n}=\{Z_{1},\ldots,Z_{n}\}$. If one takes a parametric model approach, there will be two potential dangers lurking behind the scene: (i) model misspecification (e.g., Claeskens and Hjort 2008) and (ii) the effect of selection (e.g., Leeb 2009; Berk et al. 2013; Hong et al. 2018; Kuchibhotal et al. 2022). While these two issues can be largely avoided by employing a nonparametric model, nearly all nonparametric models have tuning parameters and are only asymptotically valid. These issues associated with a model-based approach prompted researchers to seek a model-free approach for creating valid prediction regions. Early works in this direction include Wilks (1941), Fligner and Wolfe (1976), and Frey (2013). However, these works only treat the unsupervised learning case. It is unclear how the methods proposed by these papers can be applied to the regression problem without disregarding the information supplied by predictors. For prediction in the regression setting, conformal prediction (e.g., Vovk et al. 2005, 2009; Shafer and Vovk 2008; Barber et al. 2021) is a model-free machine learning method for generating finite-sample valid prediction regions at a given confidence level.
To apply conformal prediction, we first choose a non-conformity measure $M(B,z)$ which is a real-valued deterministic mapping of two arguments, where the first argument $B=\{z_{1},\ldots,z_{n}\}$ is a bag, i.e., a collection, of observed data and the second argument $z=(x,y)$ is a provisional value of a future observation. Then we run the following Algorithm 1:
@@ 5 Return $\mathsf{pl}_{x_{n+1},z^{n}}(y)$;
$C_{\alpha}(x;z^{n})=\{y:\mathsf{pl}_{x_{n+1},Z^{n}}(y)>\alpha\},$ (1)
where $0<\alpha<1$. The basic properties of the rank statistic imply the next theorem.
@@
Suppose $Z_{1},Z_{2},\ldots$ is a sequence of exchangeable random vectors and each $Z_{i}$ is generated from a distribution $\mathsf{P}$. Let $\mathsf{P}^{n+1}$ denote the corresponding joint distribution of $Z^{n+1}=\{Z_{1},\ldots,Z_{n},Z_{n+1}\}$. For $\alpha\in(0,1)$, define $t_{n}(\alpha)=(n+1)^{-1}\lfloor(n+1)\alpha\rfloor$, where $\lfloor a\rfloor$ denotes the greatest integer less than or equal to $a$. Then
@@
$\mathsf{P}^{n+1}\{Y_{n+1}\in C_{\alpha}(X_{n+1};Z^{n})\}\geq 1-\alpha\quad\text{for all $(n,\mathsf{P})$},$ (2)
@@
While (2) guarantees that the conformal prediction region $C_{\alpha}(x;Z^{n})$ is finite-sample validity for any non-conformity measure $M$, a data scientist faces several challenges in practice. First, it is clear from (1) that exact determination of a conformal prediction region generally requires one to run Algorithm 1 for all possible $y\in\mathbb{R}$. This is practically impossible. This is the reason why most existing papers only implement Algorithm 1 for a grid of $y$ values. Let $\widehat{C}_{\alpha}(Z_{n+1};Z^{n})$ denote the set results from such an approximation. Then the finite-sample validity of $\widehat{C}_{\alpha}(Z_{n+1};Z^{n})$ is nowhere justified. It is important to underline that this challenge is not to be confused with the finite-sample validity of the conformal prediction region $C_{\alpha}(X_{n+1};Z^{n})$. The finite sample validity of $C_{\alpha}(X_{n+1};Z^{n})$ is guaranteed by Theorem 1. But the set $\widehat{C}_{\alpha}(Z_{n+1};Z^{n})$ is not the same as the conformal prediction region $C_{\alpha}(X_{n+1};Z^{n})$. Therefore, one cannot use the finite-sample validity of $C_{\alpha}(X_{n+1};Z^{n})$ to justify the finite-sample validity of $\widehat{C}_{\alpha}(X_{n+1};Z^{n})$. The second challenge is closely related to the first. Even if we choose to consider only a grid of $y$ values, the computation needed for a determining $\widehat{C}_{\alpha}(Z_{n+1};Z^{n})$ could still be prohibitively expensive, although some methods, such as split conformal prediction, have been proposed to circumvent this challenge; see, for example, Lei et al. (2018). Finally, the prediction region $C_{\alpha}(X_{n+1};Z^{n})$ is not guaranteed to be an interval. In general, $C_{\alpha}(Z_{n+1};Z^{n})$ can be a disjoint union of several non-overlapping intervals, which is inappropriate for many applications; see, for example, Lei et al. (2013). For many practical purposes, a data scientist often needs a prediction region to be an interval of a certain shape, such as $(-\infty,a)$, $(a,b)$, or $(b,\infty)$, where $a$ and $b$ are real numbers.
Among these three challenges, the first one is the most serious one because it affects the finite-sample validity of prediction—the key selling point of conformal prediction. However, it is rarely addressed in the literature. Two exceptions are Hong and Martin (2021) and Hong (2025). Hong (2025) finds a suitable non-conformity measure $M$ so that
@@
where $X^{n+1}=\{X_{1},\ldots,X_{n},X_{n+1}\}$ and $f$ is some real-valued function. Then, (3) implies that the corresponding plausibility function is monotonic in $y$, which further implies the resulting conformal prediction region $C_{\alpha}(X_{n+1},Z^{n})$ equals the prediction interval based on some order statistics. The ad hoc strategy employed by Hong and Martin (2021) for unsupervised learning is of the same spirit, though their choice of their non-conformity measure leads to
@@
where $X_{i}\in\mathbb{R}$ for $1\leq i\leq n+1$. In both cases, the three aforementioned challenges are overcome simultaneously: the determination of $C_{\alpha}(X_{n+1},Z^{n})$ is exact, the computation needed is simple, and the shape of the $C_{\alpha}(X_{n+1},Z^{n})$ is an interval. In general, if we can determine $C_{\alpha}(X_{n+1},Z^{n})$ exactly, then the computational challenge is likely to be vanish, though care is still needed to ensure $C_{\alpha}(X_{n+1},Z^{n})$ is an interval of a desired form.
Recall that a non-conformity measure $M$ is said to be monotonically increasing if
$y\leq y^{\prime}\Longrightarrow M(B,y)\leq M(B,y^{\prime});$
$M$ is said to be monotonically decreasing if
$y\leq y^{\prime}\Longrightarrow M(B,y)\geq M(B,y^{\prime}).$
We say $M$ is monotonic if it is either monotonically increasing or monotonically decreasing.
Given the above observations regarding Hong and Martin (2021) and Hong (2025), it is natural to ask whether we can say something about the relationship among the monotonicity of $M$, the monotonicity of the plausibility function, Property (3), the shape of the conformal prediction region, and exact determination of the conformal prediction region $C_{\alpha}(X_{n+1},Z^{n})$. In particular, there are several open questions:
@@
  9. (IX)
In this article, we offer new insights into conformal prediction by answering the above questions. These new insights suggest that it is challenging to find a hard-and-fast rule for choosing a non-conformity measure. In view of this fact and the principle of parsimony, we propose a simple strategy to overcome the aforementioned three challenges simultaneously. The remainder of the paper is organized as follows. Section 2 answers Questions (I)–(IX). Section 3 details our general strategy to overcome the three practical challenges of conformal prediction. Section 4 provides two numerical examples based simulation to demonstrate the excellent performance of the proposed method. Section 5 concludes the article with some remarks. The insights in Section 2 and the strategy in Section 3 have their counterparts in unsupervised learning, which are given in the Appendix.

### v2

@@ Abstract
Conformal prediction is a model-free machine learning method for constructing prediction regions at a guaranteed coverage probability level. However, a data scientist often faces three challenges in practice: (i) the determination of a conformal prediction region is only approximate, jeopardizing the finite-sample validity of prediction, (ii) the computation required could be prohibitively expensive, and (iii) the shape of a conformal prediction region is hard to control. This article offers new insights into the relationship among the monotonicity of the non-conformity measure, the monotonicity of the plausibility function, and the exact determination of a conformal prediction region. Based on these new insights, we propose a quadratic-polynomial non-conformity measure that allows a data scientist to circumvent the three challenges simultaneously within the full conformal prediction framework.
Keywords and phrases: Data science; exact determination of conformal prediction regions; explainable machine learning; finite-sample validity; full conformal prediction.
@@ 1 Introduction
1.1 Conformal prediction and three practical challenges
Consider the typical regression setting where the true data-generating mechanism is
$Y=f(X)+\varepsilon,$ (1)
where $Y\in\mathbb{R}$ is the response variable, $X\in\mathbb{R}^{p}$ for some $p\geq 1$ is the vector of $p$ predictors, and $\varepsilon$ is the random error term. It is customary to assume $\mathsf{E}[\varepsilon]=0$. We do not adopt this assumption, since it is unnecessary for the method proposed in this article. Suppose $Z_{1}=(X_{1},Y_{1}),Z_{2}=(X_{2},Y_{2}),\ldots$ is a sequence of exchangeable observations from (1), where each $Z_{i}$ follows a distribution $\mathsf{P}$. Our goal is to perform an interval prediction of the next response $Y_{n+1}$ at a randomly sampled feature $X_{n+1}$, based on past observations of $Z^{n}=\{Z_{1},\ldots,Z_{n}\}$. If one takes a parametric model approach, there will be two potential dangers lurking behind the scene: (i) model misspecification (e.g., Claeskens and Hjort 2008) and (ii) the effect of selection (e.g., Leeb 2009; Berk et al. 2013; Hong et al. 2018; Kuchibhotal et al. 2022). While these two issues can be largely avoided by employing a non-parametric model, predictions based on a non-parametric model are usually only asymptotically valid. These issues associated with a model-based approach prompted researchers to seek a model-free approach for creating valid prediction regions. Early works in this direction include Wilks (1941) and Fligner and Wolfe (1976). However, these authors only treated the unsupervised learning case, and their methods are not immediately applicable in the regression setting. For prediction in the regression setting, conformal prediction (e.g., Vovk et al. 2005, 2009; Shafer and Vovk 2008; Barber et al. 2021) is a model-free machine learning method for generating finite-sample valid prediction regions at a given confidence level.
To apply conformal prediction, we first choose a non-conformity measure $M(B,z)$ which is a real-valued deterministic mapping of two arguments, where the first argument $B=\{z_{1},\ldots,z_{n}\}$ is a bag, i.e., a collection, of observed data and the second argument $z=(x,y)$ is a provisional value of a future observation $Z_{n+1}$. Then we run the following Algorithm 1:
@@ 4 Compute $\mathsf{pl}_{x_{n+1},z^{n}}(y)=(n+1)^{-1}\sum_{i=1}^{n+1}1\{\mu_{i}\geq\mu_{n+1}\}$;
$C_{\alpha}(X_{n+1},Z^{n})=\{y:\mathsf{pl}_{X_{n+1},Z^{n}}(y)>t_{n}(\alpha)\},$ (2)
where $0<\alpha<1$, $t_{n}(\alpha)=(n+1)^{-1}\lfloor(n+1)\alpha\rfloor$, and $\lfloor a\rfloor$ denotes the greatest integer less than or equal to $a$. The basic properties of the rank statistic imply the next theorem.
@@
Suppose $Z_{1},Z_{2},\ldots$ is a sequence of exchangeable random vectors and each $Z_{i}$ is generated from a distribution $\mathsf{P}$. Let $\mathsf{P}^{n+1}$ denote the corresponding joint distribution of $Z^{n+1}=\{Z_{1},\ldots,Z_{n},Z_{n+1}\}$. For $\alpha\in(0,1)$, define $t_{n}(\alpha)=(n+1)^{-1}\lfloor(n+1)\alpha\rfloor$. Then
@@
$\mathsf{P}^{n+1}\{Y_{n+1}\in C_{\alpha}(X_{n+1},Z^{n})\}\geq 1-\alpha\quad\text{for all $(n,\mathsf{P})$},$ (3)
@@
While (3) guarantees that the conformal prediction region $C_{\alpha}(X_{n+1},Z^{n})$ is finite-sample valid for any non-conformity measure $M$, a data scientist faces several challenges in practice. First, it is clear from (2) that exact determination of a conformal prediction region generally requires a data scientist to run Algorithm 1 for all possible $y\in\mathbb{R}$. This is practically impossible since there are infinitely many possible $y$-values. To be clear, by exact determination we mean the process of determining $C_{\alpha}(X_{n+1},Z^{n})$ must not involve any approximation. Clearly, a brute-force approach for running Algorithm 1 can only be done for a grid of $y$ values. Let $\widehat{C}_{\alpha}(X_{n+1},Z^{n})$ denote the prediction set results from a brute-force approach. Then the finite-sample validity of $\widehat{C}_{\alpha}(X_{n+1},Z^{n})$ is no longer guaranteed by (3). It is important to underline that this challenge is not to be confused with the finite-sample validity of the conformal prediction region $C_{\alpha}(X_{n+1},Z^{n})$. The finite sample validity of $C_{\alpha}(X_{n+1},Z^{n})$ is guaranteed by Theorem 1. But the set $\widehat{C}_{\alpha}(X_{n+1},Z^{n})$ is not the same as the conformal prediction region $C_{\alpha}(X_{n+1},Z^{n})$. Therefore, one cannot use the finite-sample validity of $C_{\alpha}(X_{n+1},Z^{n})$ to justify the finite-sample validity of $\widehat{C}_{\alpha}(X_{n+1},Z^{n})$. The second challenge is closely related to the first. Even if we choose to consider only a grid of $y$ values, the computation needed for determining $\widehat{C}_{\alpha}(X_{n+1},Z^{n})$ could still be prohibitively expensive. Finally, the prediction region $C_{\alpha}(X_{n+1},Z^{n})$ is not guaranteed to be an interval. In general, $C_{\alpha}(X_{n+1},Z^{n})$ can be a disjoint union of several non-overlapping intervals (e.g., Lei et al. 2013), which is inappropriate for many applications. For practical purposes, a data scientist often needs a prediction region to be an interval of a certain shape, such as $(-\infty,a)$, (b,b)(-b,b), or $(c,\infty)$, where $a,b$, and $c$ are real numbers and $b>0$. We will only aim at these three shapes for $C_{\alpha}(X_{n+1},Z^{n})$ in this article. Among these challenges, the first is the most serious one, because it can compromise the finite-sample validity of predictions—the key selling point of conformal prediction.
1.2 Extant approaches to addressing practical challenges of conformal prediction
In the existing literature, there are four major approaches to addressing the above challenges. Here we review these approaches in chronological order.
The first approach is to seek an easy-to-compute equivalent form of the conformal prediction set $C_{\alpha}(X_{n+1},Z^{n})$. This approach was pioneered by Vovk et al. (2005) and reviewed in Burnaev and Vovk (2014). These authors used the absolute ridge regression residual as a non-conformity measure and showed that the resulting conformal prediction set $C_{\alpha}(X_{n+1},Z^{n})$ has an equivalent form that be determined exactly (e.g., Vovk et al, Section 2.3; Burnaev and Vovk 2014). One drawback of their proposal is that their conformal prediction set $C_{\alpha}(X_{n+1},Z^{n})$ is not necessarily an interval, and the sufficient condition for their conformal prediction set to be an interval is data-dependent (e.g., Vovk et al., Pages 31–32) and beyond the control of a data scientist. Lei (2019) applied linear homotopy, a tool in algebraic topology, to find equivalent set of $C_{\alpha}(X_{n+1},Z^{n})$ when the non-conformity measure is absolute lasso residual, and showed that $C_{\alpha}(X_{n+1},Z^{n})$ in this case is not necessarily an interval. He also established that $C_{\alpha}(X_{n+1},Z^{n})$ would be an interval if we replace the absolute lasso residual with the absolute elastic net residual. Li (2024) extended these results to a general non-conformity measure. However, no closed-form formula for lasso or elastic net solution is known. This means a practitioner must resort to numerical approximation to implement the “conformal lasso algorithm” in Lei (2019), and the finite-sample validity of the resulting prediction set is no longer guaranteed. The same can be said for the “fast conformalization algorithm” in Li (2024), since we generally need numerical approximation in solving ODEs (ordinary differential equations). Therefore, while Lei (2019) and Li (2024) made important contributions by chartering a new course towards the exact determination of $C_{\alpha}(X_{n+1},Z^{n})$, further research is required to either (i) obtain a closed-form formula for the solutions to lasso, elastic net, and ODEs, or (ii) show why approximation used in numerical solution to lasso, elastic net, and ODEs does not affect the finite-sample validity of the resulting prediction set. In unsupervised learning, Hong and Martin (2021) showed that if we take the non-conformity measure to be the empirical CDF (cumulative distribution function) and the kernel of the empirical CDF is constant, then the resulting conformal prediction set coincides with the one-sided prediction interval based on order statistics (e.g., Wilks 1941; Fligner Wolfe 1976) and can be determined exactly. Under the constraint that $Y$ is non-negative, Hong (2026a) showed that for a specific non-conformity measure, $C_{\alpha}(X_{n+1},Z^{n})$ can be determined exactly and is a one-sided interval.
The second approach is initiated by Lei et al. (2013, 2014). In this approach, we obtain a sandwich approximation set $C^{+}_{\alpha}(X_{n+1},Z^{n})$ of $C_{\alpha}(X_{n+1},Z^{n})$ such that (i) $C_{\alpha}(X_{n+1},Z^{n})\subseteq C^{+}_{\alpha}(X_{n+1},Z^{n})$, and (ii) $C^{+}_{\alpha}(X_{n+1},Z^{n})$ can be determined exactly. Since $C_{\alpha}(X_{n+1},Z^{n})$ is finite-sample valid, so is $C^{+}_{\alpha}(X_{n+1},Z^{n})$. One potential drawback of this approach is that we do not know how conservative $C^{+}_{\alpha}(X_{n+1},Z^{n})$ is relative to $C_{\alpha}(X_{n+1},Z^{n})$ especially when the sample size is small or modest. Lei et al (2013, 2014) obtained some convergence rate results. However, since every convergence rate result is up to an unknown constant, these results do not tell us the precise degree of conservativeness of $C^{+}_{\alpha}(X_{n+1},Z^{n})$. In addition, kernel density estimators are used to create the non-conformity measures in Lei et al. (2013, 2014). It is unclear to find a superset of $C^{+}_{\alpha}(X_{n+1},Z^{n})$ for an arbitrary non-conformity measure.
The third approach, called split conformal prediction, was proposed by Lei et al. (2018), though similar ideas appeared under the name of inductive conformal inference in Papadopoulos et al. (2002) and Vovk et tal. (2005). In this approach, we first randomly split the data into two equal-sized parts: a training set and a test set; then, we fit a regression model, such as ridge regression or lasso using the training data set; next, we apply the prediction interval based on order statistics (e.g., Wilks 1941; Fligner Wolfe 1976) to the absolute test predictive errors to obtain a finite-sample valid $100(1-\alpha)\%$ prediction interval $J_{\alpha}(X_{n+1},Z^{n})$ for the next absolute test predictive error; finally, we converts $J_{\alpha}(X_{n+1},Z^{n})$ to a finite-sample valid $100(1-\alpha)\%$ prediction set for the response variable $Y$. Split conformal prediction is widely applicable since the data scientist is free to choose a regression model, and a split conformal prediction set is guaranteed to be a finite-sample valid interval. For example, if we replace the predicted value of $Y$ by the value of the quantile function at $X_{i}$ in the split conformal prediction, we will obtain the quantile regression conformal prediction set proposed in Romano et al. (2019). However, if the chosen regression model is wrong, each test absolute predictive error will likely be large, rendering the resulting split conformal prediction interval conservative. To avoid confusion, we will refer to the original conformal prediction framework as full conformal prediction. When no confusion might arise, we might suppress the word “full”. Therefore, “conformal prediction” means “full conformal prediction”, unless otherwise stated.
The fourth approach is based on a general strategy advanced by Hong (2026b). In this approach, we first choose a non-constant $p$-variate real-valued function $h$, called a transformation, and write (1) as
$Y=h(X)+W,$
where W=Yh(X)W=Y-h(X). Then, we obtain $W_{i}=Y_{i}-h(X_{i}),\ i=1,\ldots,n$ using data $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$. Next, we obtain a finite-sample valid $100(1-\alpha)\%$ prediction interval $I_{\alpha}(X_{n+1},Z^{n})$ for $W_{n+1}$, using order statistics (e.g., Wilks 1941; Fligner Wolfe 1976). Finally, we transfer $I_{\alpha}(X_{n+1},Z^{n})$ to a finite-sample valid $100(1-\alpha)\%$ for $Y_{n+1}$ by a shift of $h(X_{n+1})$. We will call this approach transformation method. Similar to split conformal prediction, the transformation method is also widely applicable, since a data scientist can choose any transformation $h$, and the resulting prediction set is always a finite-sample valid interval. However, no theoretical result is available for choosing the transformation $h$, though empirical evidence in Hong (2026b) suggests that if $h$ grows no faster than the linear function in each of its variables, the resulting prediction interval will be relatively tight.
Note that only the first two approaches are within the full conformal prediction framework. Both split conformal prediction and the transformation method have moved away from full conformal prediction. In particular, no leave-one-out action is performed to compute a non-conformity score in these two methods.
1.3 Our contributions
This article follows the first approach to investigate how we might overcome the above challenges in the full conformal prediction framework. Following Vovk et al. (2019), we say a non-conformity measure $M$ is monotonically increasing if
$y\leq y^{\prime}\Longrightarrow M(B,(x,y))\leq M(B,(x,y^{\prime}))\quad\text{for all $x\in\mathbb{R}^{p}$};$
we say $M$ is monotonically decreasing if
$y\leq y^{\prime}\Longrightarrow M(B,(x,y))\geq M(B,(x,y^{\prime}))\quad\text{for all $x\in\mathbb{R}^{p}$}.$
We say $M$ is monotonic if it is either monotonically increasing or monotonically decreasing. Monotonicity of a non-conformity in unsupervised learning is defined similarly; see the Appendix. Hong (2026a) found a suitable monotonic non-conformity measure $M$ so that
@@ 1.3 Our contributions
where $X^{n+1}=\{X_{1},\ldots,X_{n},X_{n+1}\}$ and $f$ is some real-valued function. Then, (4) implies that the corresponding plausibility function is monotonic in $y$, which further implies the resulting conformal prediction region $C_{\alpha}(X_{n+1},Z^{n})$ equals the prediction interval based on some order statistics. The ad hoc strategy employed by Hong and Martin (2021) for unsupervised learning is of the same spirit; their non-conformity measure is also monotonic, and their non-conformity measure leads to
@@
where $X_{i}\in\mathbb{R}$ for $1\leq i\leq n+1$. In both cases, the three aforementioned challenges are overcome simultaneously: the determination of $C_{\alpha}(X_{n+1},Z^{n})$ is exact, the computation needed is simple, and the shape of the $C_{\alpha}(X_{n+1},Z^{n})$ is an interval. In general, if we can determine $C_{\alpha}(X_{n+1},Z^{n})$ exactly, then the computational challenge will likely vanish, though care is still needed to ensure $C_{\alpha}(X_{n+1},Z^{n})$ is an interval of a desired form.
Given the above observations regarding Hong and Martin (2021) and Hong (2026a), it is natural to ask whether we can say something about the relationship among the monotonicity of $M$, the monotonicity of the plausibility function, Property (4), the shape of the conformal prediction region, and exact determination of the conformal prediction region $C_{\alpha}(X_{n+1},Z^{n})$. In particular, there are two sets of open questions.
Questions regarding monotonicity of $M$
@@
Questions regarding general properties of full conformal prediction
  1. (VI)
@@
The contributions of this article are multi-fold. First, we offer new insights into full conformal prediction by answering the above questions. These new insights suggest that it is challenging to find a hard-and-fast rule for choosing a non-conformity measure. In view of this fact and the principle of parsimony, we then propose a flexible quadratic polynomial non-conformity measure. We show that by setting the parameters of this quadratic polynomial non-conformity measure to certain values, the corresponding $100(1-\alpha)\%$ conformal prediction interval $C_{\alpha}(X_{n+1},Z^{n})$ is a one-sided interval or two-sided interval, and these intervals can be determined easily and exactly. In particular, the (inflexible) linear non-conformity measure proposed in Hong (2026a) is a special case of our (flexible) quadratic polynomial non-conformity measure. Therefore, our proposed non-conformity measure is one of the rare ones, if not the first one, in the full conformal prediction framework that leads to exact determination of $C_{\alpha}(X_{n+1},Z^{n})$, and allows a data scientist to choose whether $C_{\alpha}(X_{n+1},Z^{n})$ is $(-\infty,a)$-shaped, $(a,\infty)$-shaped, or $(a,b)$-shaped. In addition, we establish an interesting result regarding the efficiency of a bounded finite-sample prediction interval: there is no bounded finite-sample valid non-parametric $100(1-\alpha)\%$ prediction interval (based on conformal prediction or not) with the shortest mean length for any confidence level $1-\alpha$.
The remainder of the paper is organized as follows. Section 2 answers Questions (I)–(IX). Section 3 details our proposal. Section 4 provides several numerical examples, based on simulation, to demonstrate the performance of the proposed method. Section 5 concludes the article with some remarks. The insights in Section 2 and the strategy in Section 3 have their counterparts in unsupervised learning, which are given in the Appendix.